# Very Basic Set theory

### v4

@@
Therefore Very basic set theory is a set theory with individuals.
1. What are individuals?
@@
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is ‘contained’ in $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that is for every $s$ and for every $t$ either $s\in t$ or $s\notin t$ but not both. If $s\in s$ then $s$ is an individual.
We prove Russell’s paradox
@@
For every $s$ and for every $t$ the membership relation $s\in t$ is definite. Thus
if $s\in s$ then exists $u$ such that ($u\in s$ and $u\in u$),
if $s\notin s$ then exists $u$ such that ($u\notin s$ and $u\notin u$). Then for every $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$.
Zermelo’s 1908 axiomatization permitted individuals. And the prove of Russell’s paradox includes individuals and other sets. But a set which contains itself, an individual, is problematic. (Boolos [1971] p.217, pp.219-220)
@@
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. (Boolos [1971] p.215)
”Cantor also defined a set as a ‘many, which can be thought of as one, i.e. a
totality of definite elements that can be combined into a whole by a law.’ ”
Then ”which can be thought of as one,” thus a set is a ‘mental picture’ in your mind. Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Logic treats every possibility, and $s\in s$ is possible. Then someone reading or listening that language forms a mental picture of a set in her/his mind. The picture in her/his mind agrees or disagrees with the picture in your mind. We call that ‘experience’.
What is an individual? ‘Set $s$ is an individual’ is what is called in (Wittgenstein [1999]) ”an atomic fact”. In the Introduction (p.12) Bertrand Russell explains
@@ 2. What is a set?
Set theory follows rules of logic but the membership relation is definite. That means the logic of ‘true’ and ‘false’ should be changed to the logic of ‘definiteness’.
In the next section we prove logical formulas preserve definiteness.
@@ 3. Logic and definiteness
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise exists $s$ such that $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ and $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $(\Phi\wedge\neg\Phi)$.
@@
There are many rewite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
@@ 9. Axiom ‘union’
The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and axiom regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘natural numbers’.
@@ Conclusion
Very basic set theory is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first. But individuals are ‘not needed for mathematical purposes’ (Oliver and Smliley [2006] §5.1). Thus Russell’s paradox is not the result of the logic of definiteness otherwise individuals exist, and the naming of a special set is logically first (an ‘empty set’, or an ‘infinite set’). Individuals are banished from ‘mainstream set theory’.
There are more differences with ‘mainstream set theory’. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.

### v5

@@
Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
Introduction
@@
Sets are denoted by $s$, $t$, $u$, …. Symbol $\in$ is the ‘membership relation’ symbol. Then $s\in t$ means $s$ is a ‘constituent’ (also called a member) of set $t$. The negation of $s\in t$ is denoted by $s\notin t$. The membership relation is definite, that means
  • for every $s$ and for every $t$ applies either $s\in t$ or $s\notin t$, but not both.
Then for every $s$ it is the case $s\in s$ or $\notin s$. We prove Russell’s paradox
@@
1. What is an individual?
In set theory $s\in s$ is possible thus individuals are possible. But a set that contains itself, an individual, is problematic (Boolos [1971] p.217, pp.219-220)
@@ 1. What is an individual?
However, ‘binding up some things’ or ‘containing something’ is not the definition of a set. A set does not ‘contain’ something, the constituents of a set exist.
Georg Cantor defined a set as (Boolos [1971] p.215)
  • ”a ‘many, which can be thought of as one, i.e. a totality of definite
    elements that can be combined into a whole by a law’.”
Thus a set is something which can be thought of as one, therefore a set is a ‘mental picture’ in your mind. That is in line with (Wittgenstein [1999] p.39) proposition
  • ”2.1 We make to ourselves pictures of facts.”
Therefore we prefer to write ‘constituent of a set’, not ‘contained in a set’.
Set theory is a language that follows rules of logic to communicate the mental picture of a set with the physical world. Then someone reading, seeing or listening that language forms a mental picture of a set. That picture agrees or disagrees with the picture in your mind. (Wittgenstein [1999] p.43) proposition
  • ”2.222 In the agreement or disagreement of its sense with reality, its truth
    or falsity consists.”
An individual is what is called in (Wittgenstein [1999]) ”a simple sign” (propositions
3.2, 3.201, 3.202, 3.26). In the Introduction (p.12) Bertrand Russell explains
@@ 2. What is a set?
Set theory follows rules of logic. Logic treats every possibility thus logic treats the definiteness of the membership relation. In the next section we prove the building up of logical formulas preserves definiteness.
@@ 3. Logic and definiteness
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
@@
There are many rewrite rules: ¬($\neg(\neg\Phi)$ is logically equivalent to $\Phi$, $\neg(\Phi\leftrightarrow\Psi)$ is logically equivalent to $(\Phi\leftrightarrow\neg\Psi)$ and $((\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi))$ is logically equivalent to $((\Phi\wedge\Psi)\wedge(\neg\Phi\vee\neg\Psi))$. More rewrite rules in (Mendelson [2015] p.15).
@@ 9. Axiom ‘union’
The definition of equality and axiom pairs are only meaningful if exist sets the naming of which are logically first. Thus with individuals and axiom regularity. And regularity is used with axiom union to prove that for every set the transitive closure exists. The proof needs no other axioms and no ‘ set of all natural numbers’.
@@ Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first. But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.
There are more differences with mainstream set theory. For example, the axiom ‘powerset’ is as usual but Cantor’s Theorem does not hold.